# OneMall: One Model, More Scenarios — End-to-End Generative Recommender Family at Kuaishou E-Commerce

### v1

@@ Abstract.
In the wave of generative recommendation, we present OneMall, an end-to-end generative recommendation framework tailored for e-commerce services at Kuaishou. Our OneMall systematically unifies the e-commerce’s multiple item distribution scenarios, such as Product-card, short-video and live-streaming. Specifically, it comprises three key components, aligning the entire model training pipeline to the LLM’s pre-training/post-training: (1) E-commerce Semantic Tokenizer: we provide a tokenizer solution that captures both real-world semantics and business-specific item relations across different scenarios; (2) Transformer-based Architecture: we largely utilize Transformer as our model backbone, e.g., employing Query-Former for long sequence compression, Cross-Attention for multi-behavior sequence fusion, and Sparse MoE for scalable auto-regressive generation; (3) Reinforcement Learning Pipeline: we further connect retrieval and ranking models via RL, enabling the ranking model to serve as a reward signal for end-to-end policy retrieval model optimization. Extensive experiments demonstrate that OneMall achieves consistent improvements across all e-commerce scenarios: +13.01% GMV in product-card, +15.32% Orders in Short-Video, and +2.78% Orders in Live-Streaming. OneMall has been deployed, serving over 400 million daily active users at Kuaishou.
@@ 1. Introduction
With advances in pre-training, post-training, and infrastructure, the bottleneck of training on tons of Tokens at hundreds of Billion parameters within acceptable time and computation resource limitations has been significantly reduced and possible, enabling the LLM community to continuously incubate advanced artificial intelligence toward human-level performance, rise to many superintelligent systems such as Chat-Bot (Team et al., 2025a), Vision-Generator (Wu et al., 2025), Code-Agent (Team et al., 2025b) and others. Given such success of LLMs, many industrial researchers have been pondering one thing: can we replicate the LLM evolution trace at pre-training/post-training in RecSys?
@@
  • We conduct extensive offline and online experiments at Kuaishou e-commerce service. The offline experiments show that all prediction tasks get significant improvements, and the online experiments +13.01% GMV in product-card, +15.32% Orders in Short-Video, and +2.78% Orders in Live-Streaming.
@@ 4.3. Online Experiments
From the results, OneMall achieves significant improvements across all scenarios: (1) In Product-Card, OneMall obtains +2.851% Exposure, +13.01% Click, +8.35% Order, and +14.71% GMV, demonstrating strong commercial value in the Mall Tab where users have explicit purchase intent. (2) In Short-Video, we observe +15.32% Exposure and +11.65% Order improvements, indicating that our tokenizer effectively captures both watching experience and commercial intent for e-commerce short-videos. (3) In Live-Streaming, OneMall achieves +4.47% Order and +4.90% GMV improvements, validating the effectiveness of our dynamic Semantic ID generation strategy for live-streaming content. Overall, these online results confirm that OneMall successfully bridges the gap between user interest modeling and commercial conversion, bringing substantial business value to Kuaishou’s e-commerce platform.
@@ 4.4. RL Effectiveness Analysis
In this section, we thoroughly validate the effectiveness of reinforcement learning in e-commerce recommendation. The core business objectives in e-commerce are CTR, CTCVR, and EGPM. We first perform score fusion as defined in Eq.(8), with $\alpha=1.0$, $\beta=30.0$, $\gamma=1.0$ to maintain consistent magnitude across different objectives. We sample 2% of training traffic for RL training with a beam size of 768 for candidate generation. Additionally, we set the loss weight ratio ($\mathcal{L}_{\texttt{NTP}}+0.5\mathcal{L}_{\texttt{RL}}$). We empirically find that excessively large RL loss weights degrade the SID accuracy metrics and affect the validity rate of generated candidates; thus we balance generation efficiency and quality with the hand-craft weight. From Table 4, we have: (1) DPO/GRPO vs. Base: Both DPO and GRPO achieve consistent improvements over the Base model across all candidate segments (Top10/100/500). The reward scores (CTR/CTCVR/GPM) also improve at different degrees, demonstrating the effectiveness of RL in generating high-quality candidates and its capability to optimize multiple objectives simultaneously in recommendation. (2) GRPO vs. DPO: GRPO outperforms DPO across all candidate segments, with particularly notable improvements in Top10 (e.g., +0.040% CTR, +0.012% CTCVR, +0.228% GPM over DPO). The reason might be that GRPO normalizes rewards across all 768 sampled candidates to compute advantage scores, enabling the model to learn the quality distribution within each group. In contrast, DPO only uses pairwise positive-negative samples, providing less comprehensive training feedback.
@@ 4.5. Component Analyses
To address this issue, we introduce FSQ (Finite Scalar Quantization) at the last quantization layer. FSQ fixes the cluster centers in advance, ensuring uniform distribution of codes. Although this sacrifices some semantic information, it makes the partitioned space more regular. As shown in Table 5, ResKmeansFSQ combines the advantages of both approaches: it preserves semantic information from the first two Res-Kmeans layers while ensuring uniform cluster center distribution through FSQ. Auxiliary Contrastive Loss. Since Semantic IDs are relatively coarse-grained and lack item-specific fine-grained features, we introduce additional item-side information into the model. Specifically, we feed important item attributes (e.g., product category, price, shop information) into an item tower and apply in-batch contrastive learning as an auxiliary objective. Our experiments show that this auxiliary loss yields consistent improvements in hit rate (+1.5%/+1.7%/+1.7% on HR@50/100/500), demonstrating the effectiveness of incorporating fine-grained item features to complement the coarse-grained Semantic ID representation. Query-Former Compression. To avoid the higher computation cost by long sequence, we introduce the Q-former for early sequence compression. Fortunately, we find that its performance is on par with the longer sequence version, as shown in Figure 6, the GFLOPs (calculation cost per sample) is largely reduced with only minor performance degradation.

### v2

# OneMall: One Architecture, More Scenarios — End-to-End Generative Recommender Family at Kuaishou E-Commerce

@@ Abstract.
In the wave of generative recommendation, we present OneMall, an end-to-end generative recommendation framework tailored for e-commerce services at Kuaishou. Our OneMall systematically unifies the e-commerce’s multiple item distribution scenarios, such as Product-card, short-video and live-streaming. Specifically, it comprises three key components, aligning the entire model training pipeline to the LLM’s pre-training/post-training: (1) E-commerce Semantic Tokenizer: we provide a tokenizer solution that captures both real-world semantics and business-specific item relations across different scenarios; (2) Transformer-based Architecture: we largely utilize Transformer as our model backbone, e.g., employing Query-Former for long sequence compression, Cross-Attention for multi-behavior sequence fusion, and Sparse MoE for scalable auto-regressive generation; (3) Reinforcement Learning Pipeline: we further connect retrieval and ranking models via RL, enabling the ranking model to serve as a reward signal for end-to-end policy retrieval model optimization. Extensive experiments demonstrate that OneMall achieves consistent improvements across all e-commerce scenarios: +14.7% GMV in Product-Card, +10.3% GMV in Short-Video, and +4.9% GMV in Live-Streaming. OneMall has been deployed, serving over 400 million daily active users at Kuaishou.
@@ 1. Introduction
With advances in pre-training, post-training, and infrastructure, the bottleneck of training on tons of Tokens at hundreds of Billion parameters within acceptable time and computation resource limitations has been significantly reduced and possible, enabling the LLM community to continuously incubate advanced artificial intelligence toward human-level performance, rise to many superintelligent systems such as Chat-Bot (Team et al., 2025a), Vision-Generator (Wu et al., 2025), Code-Agent (Team et al., 2025b) and others. Given such success of LLMs, many industrial researchers have been pondering one thing: can we replicate the LLM evolution trace at pre-/post-training in RecSys?
@@
  • We conduct extensive offline and online experiments at Kuaishou e-commerce service. The offline experiments show that all prediction tasks get significant improvements, and the online experiments +14.7% GMV in Product-Card, +10.3% GMV in Short-Video, and +4.9% GMV in Live-Streaming.
@@ 4.3. Online Experiments
From the results, OneMall achieves significant improvements across all scenarios: (1) In Product-Card, OneMall obtains +2.851% Exposure, +13.01% Click, +8.35% Order, and +14.71% GMV, demonstrating strong commercial value in the Mall Tab where users have explicit purchase intent. (2) In Short-Video, we observe +5.76% Click and +11.65% Order and 10.33% GMV improvements, indicating that our tokenizer effectively captures both watching experience and commercial intent for e-commerce short-videos. (3) In Live-Streaming, OneMall achieves +4.47% Order and +4.90% GMV improvements, validating the effectiveness of our dynamic Semantic ID generation strategy for live-streaming content. Overall, these online results confirm that OneMall successfully bridges the gap between user interest modeling and commercial conversion, bringing substantial business value to Kuaishou’s e-commerce platform.
@@ 4.4. RL Effectiveness Analysis
In this section, we thoroughly validate the effectiveness of reinforcement learning in e-commerce recommendation. The core business objectives in e-commerce are CTR, CTCVR, and EGPM. We first perform score fusion as defined in Eq.(8), with $\alpha=1.0$, $\beta=30.0$, $\gamma=1.0$ to maintain consistent magnitude across different objectives. We sample 2% of training traffic for RL training with a beam size of 768 for candidate generation. Additionally, we set the loss weight ratio ($\mathcal{L}_{\texttt{NTP}}+0.5\mathcal{L}_{\texttt{RL}}$). We empirically find that excessively large RL loss weights degrade the SID accuracy metrics and affect the validity rate of generated candidates; thus we balance generation efficiency and quality with the hand-craft weight. From Table 4, we have: (1) DPO/GRPO vs. Base: Both DPO and GRPO achieve consistent improvements over the Base model across all candidate segments (Top10/100/500). The reward term and score also improve at different degrees, demonstrating the effectiveness of RL in generating high-quality candidates and its capability to optimize multiple objectives simultaneously in recommendation. (2) GRPO vs. DPO: GRPO outperforms DPO across all candidate segments, with particularly notable improvements in Top10 (e.g., +0.040% CTR, +0.012% CTCVR, +0.228% GPM over DPO). The reason might be that GRPO normalizes rewards across all 768 sampled candidates to compute advantage scores, enabling the model to learn the quality distribution within each group. In contrast, DPO only uses pairwise positive-negative samples, providing less training feedback.
@@ 4.5. Component Analyses
To address this issue, we introduce FSQ (Finite Scalar Quantization) at the last quantization layer. FSQ fixes the cluster centers in advance, ensuring uniform distribution of codes. Although this sacrifices some semantic information, it makes the partitioned space more regular. As shown in Table 5, ResKmeansFSQ combines the advantages of both approaches: it preserves semantic information from the first two Res-Kmeans layers while ensuring uniform cluster center distribution through FSQ. Auxiliary Contrastive Loss. Since Semantic IDs are relatively coarse-grained and lack item-specific fine-grained features, we introduce additional item-side information into the model. Specifically, we feed important item attributes (e.g., product category, price, shop information) into an item tower and apply in-batch contrastive learning as an auxiliary objective. Our experiments show that this auxiliary loss yields consistent improvements in hit rate (+1.5%/+1.7%/+1.7% on HR@50/100/500), demonstrating the effectiveness of incorporating fine-grained item features to complement the coarse-grained Semantic ID representation. Query-Former Compression. To avoid the higher computation cost by long sequence, we introduce the Q-former for early sequence compression. Fortunately, we find that its performance is on par with the longer sequence version, as shown in Table 6, the GFLOPs (calculation cost per sample) is largely reduced with only minor performance degradation.